### v1

# In-Context Reverse Classification Accuracy: Efficient Estimation of Segmentation Quality without Ground-Truth

Abstract
Assessing the quality of automatic image segmentation is crucial in clinical practice, but often very challenging due to the limited availability of ground truth annotations. In this paper, we introduce In-Context Reverse Classification Accuracy (In-Context RCA), a novel framework for automatically estimating segmentation quality in the absence of ground-truth annotations. By leveraging recent in-context learning segmentation models and incorporating retrieval-augmentation techniques to select the most relevant reference images, our approach enables efficient quality estimation with minimal reference data. Validated across diverse medical imaging modalities, our method demonstrates robust performance and computational efficiency, offering a promising solution for automated quality control in clinical workflows, where fast and reliable segmentation assessment is essential. The code is available at https://github.com/mcosarinsky/In-Context-RCA.
1 Introduction
Image segmentation plays a crucial role in many medical practices, particularly for treatment planning, clinical decision making and monitoring disease progression. In these scenarios, evaluating the performance of segmentation methods becomes of utmost importance. Typically, the evaluation is performed on annotated databases using metrics that measure the agreement between predicted segmentations and ground truth (GT) labels. Commonly used metrics include the Dice-Sørensen coefficient (DSC) and other overlap-based measures [6], as well as distance-based metrics, such as the Hausdorff distance or the average symmetric surface distance (ASSD) [16, 30].
Once a segmentation method is deployed in practice, continuously assessing its real performance on new unseen data samples becomes a challenging task, mainly due to the limited availability of reference annotations in clinical settings. However, evaluation remains a critical task, especially for identifying instances where an image segmentation method may fail in automatic pipelines due to unexpected changes, e.g. in patient demographics[10] or acquisition protocols. Here we aim to design a method that can quickly and automatically assess the quality of segmentations when there is no GT available to compare with.
We propose a flexible framework based on In-Context segmentation [32] and Reverse Classification Accuracy (RCA) [31] that we name In-Context RCA. RCA is a framework originally designed to predict the performance of an image segmentation method in scenarios where GT masks are unavailable. It works by using the predicted segmentation of a new image as a pseudo-ground truth to train a reverse classifier (a segmentation network in our case) [^1] Note that we use the term reverse classifier as introduced in the original publication [31], although it would be more accurately described as a reverse segmenter. We will use both terms interchangeably.. This model is then applied to a set of reference images with known GT, and its performance is evaluated using standard metrics like the DSC. Here we propose to combine this framework with In-Context Learning (ICL), as it naturally handles such few-shot scenarios, proposing to eliminate the need of tranining a reverse segmenter model. ICL is a paradigm that allows models to adapt to new tasks by conditioning on a small set of labeled examples, without requiring additional training. For image segmentation, it means that we can segment a target image by providing a set of example images with their corresponding segmentation masks at inference time, without re-training our segmenter [4, 37]. By leveraging in-context segmentation models as reverse classifiers, our method facilitates segmentation quality estimation at the individual image level.
Related work. Despite its relevance, few approaches have addressed automatic quality control of segmentations thus far. Many previous works have relied on extracting features from segmentations and fitting a model to these features. Such approaches have been applied to prostate segmentation [29], cardiac segmentation [33], and a more generic framework has been used to assess the quality for automated segmentations for lung and liver in [13]. However, these methods are constrained by the feature extraction, which can limit their adaptability and generalizability to diverse clinical data. Additionally, a CNN-based approach was developed for automated quality control of left ventricle segmentation in cardiac MRI [9]. In spite of demonstrating the potential of deep learning for automatic quality assessment, this method requires generating a synthetic dataset and further training multiple CNN models.
Other previous studies rely on a reverse-testing strategy, where a new classifier is trained on predictions on the test data and evaluated again on the training data [8, 36]. The limitation of these approaches is that they only provide an estimate of the average performance rather than case-specific estimations. Building on this idea, Reverse Classification Accuracy (RCA) has been shown to provide accurate predictions of segmentation quality on a per-image basis [31, 24]. In [31], Atlas-based segmentation has shown to be the most effective reverse classifier for medical image segmentation. However, the computational time required to complete atlas based segmentation on a single instance is a limitation, making it impractical for real-time quality control. In [23] the authors propose using a neural network to directly estimate prediction quality, aiming to mitigate the time constraint. However, this approach necessitates training a separate CNN for each task, which significantly limits its practicality and scalability. More importantly, both methods [31, 24] have only been used for anatomical segmentation and not for lesions, which are usually more challenging due to the lack of regularity.
Recently, Few-shot Semantic Segmentation (FSS) methods –which extend in-context learning to address segmentation tasks– have shown great potential in the medical domain, particularly with limited annotations [19]. While traditional approaches are often based on prototypical networks [27, 18], more advanced models such as UniverSeg [4] and foundational models like SAM 2 [22, 3, 37] delivered more promising results on a wide range of medical tasks. This enhanced generalizability is largely due to their training on vast amounts of diverse data. Here we will use FSS to improve RCA performance and efficiency.
Contributions. We propose a novel approach titled In-Context RCA, which extends the traditional RCA framework by leveraging in-context segmentation models as reverse classifiers. These models can adapt their predictions based on a limited number of labeled examples without the need for additional training, making them ideal for this setting. Furthermore, inspired by the recent success of retrieval augmented generation frameworks [14] for natural language processing, we introduce a more efficient variant of RCA entitled Retrieval Augmented RCA, which incorporates image embedding vector databases to efficiently retrieve the most relevant reference images to be used as few-shot examples, further enhancing efficiency and accuracy. We provide extensive validation using 10 medical imaging datasets, covering a wide range of image modalities and organs, broadening the scope of RCA analysis conducted so far and demonstrating the method’s capability to provide accurate estimates in a wide range of scenarios while drastically reducing the required inference time.
2 In-Context RCA
Let us say we deployed to production an X-ray lung segmentation model, and we want to monitor the quality of the segmentations $S_{I}$ produced for new images $I$. Let us also assume that we have at our disposal a database of reference images $J_{i}$ and GT segmentation masks $S_{J_{i}}^{GT}$ (in our case, X-ray images with expert annotations). The core idea behind RCA is to assess the quality of a segmentation $S_{I}$ corresponding to a new image $I$ by training a reverse segmenter $f_{I,S_{I}}$ (or reverse classifier) using only $S_{I}$ as pseudo-ground truth for training. The reverse segmenter is then applied to the reference database containing images $J_{i}$ with known GT masks $S_{J_{i}}^{GT}$, generating segmentations $f_{I,S_{I}}(J_{i})=S_{J_{i}}$. The quality of these segmentations is then assessed by computing a chosen quality metric $\rho$ (e.g DSC) for each $S_{J_{i}}$ and its corresponding ground truth $S_{J_{i}}^{GT}$. Finally, we estimate the metric $\rho$ for $S_{I}$ as the best metric value on the reference database:
$\hat{\rho}(S_{I})=\max_{k}\rho(S_{J_{k}},S_{J_{k}}^{GT}),$ (1)
assuming that higher values of $\rho$ correspond to better quality estimations. The underlying hypothesis is that if $S_{I}$ is of high quality, the RCA segmenter should perform well on at least some of the reference images, resulting in high values of $\rho$. Conversely, if $S_{I}$ is of poor quality, the segmenter is expected to perform poorly on the reference database.
Figure 1: General outline of the In-Context RCA framework. The quality of a predicted segmentation $S_{I}$ on an image $I$ is estimated by leveraging $(I,S_{I})$ as the support set of an In-Context Classifier or segmenter. This classifier is then applied to segment a reference dataset, selected through retrieval augmentation. Finally, the best segmentation result based on an evaluation metric $\rho$, is used to predict the quality of $S_{I}$ following Eq 1.
In this work, we propose to adopt in-context segmentation models as reverse segmenters, resulting in In-Context RCA (see Fig. 1). Unlike previous approaches, that rely on training CNNs [23] or computationally expensive atlas-based registration methods as reverse classifiers [31, 24], in-context learning models are computationally efficient and can adapt their predictions based on a small support set. We also introduce a dynamic selection of the reference dataset on a per-case basis based on retrieval similarity, enabling the use of smaller, more relevant reference datasets. Here we employ two different reverse in-context segmenters to demonstrate the approach’s versatility.
UniverSeg-based reverse segmenter. The UniverSeg architecture [4] was designed to tackle unseen medical segmentation tasks without requiring additional training. Given a query image $Q$ and a small support set $\mathcal{S}$ of image-label pairs that define a new segmentation task, UniverSeg is able to segment the query image $Q$ following the provided examples, without updating the neural weights. It follows a UNet-like encoder-decoder architecture composed of CrossBlock modules. These modules, which are convolutional layers, enable the transfer of relevant information between the support set and the query image, allowing the model to perform accurate segmentation without task-specific retraining. It was trained on a collection of 53 open-access medical segmentation datasets called MegaMedical, covering a wide range of image modalities and structures.
Here we adopt UniverSeg as the reverse classifier of the In-Context RCA framework. To this end, we use the target image $I$ and the predicted segmentation under assessment $S_{I}$ as the UniverSeg support set, i.e. we simply define $\mathcal{S}=\{(I,S_{I})\}$. We then segment all the reference images $J_{i}$ in our RCA reference database by considering them as query images, one at a time, and finally compare each resulting segmentation $S_{J_{k}}$ with the corresponding reference segmentation $S_{J_{k}}^{GT}$ to estimate $\hat{\rho}(S_{I})$ as defined in Eq. 1.
SAM 2-based reverse segmenter. Segment Anything Model (SAM) [12] is a foundation model for image and video segmentation guided by prompts. Its architecture consists of image and prompt encoders that extract features, and a mask decoder that fuses this information to generate segmentations. SAM 2 [22] extends its predecessor by incorporating memory components (an encoder, bank, and attention mechanism) enabling adaptation to new tasks and outperforming existing video object segmentation models, especially for temporal tracking. For medical FSS [3], SAM 2 treats the support set as initial frames of a sequence, with their segmentation masks serving as prompts. Despite having no temporal connection between images, this approach effectively transfers segmentation masks from annotated to unannotated examples (see [3] for more details). In our framework, we use SAM 2 as reverse classifier, following the few-shot setting with sequences that comprise only two frames: 1) the target image and its pseudo-ground truth, and 2) a query image drawn from the reference dataset.
Retrieval Augmented (RA)-RCA. We propose to reduce the size of the RCA reference database by dynamically retrieving the most relevant reference images on a per-case basis. This allows the model to focus on smaller, more relevant subsets of data, optimizing computational efficiency while maintaining and even improving the quality of predictions. To implement this, we embed all the images from the reference database using DINOv2 embeddings [17] and then create a vector database using FAISS [7], which enables efficient similarity search. For each incoming target image whose segmentation quality we want to assess, we compute its embedding and retrieve the $k$ most similar images from the reference index based on cosine similarity. While DINOv2 was not explicitly trained for the medical domain, it has shown strong generalization, making it effective for general tasks, including those in medical imaging. In fact, a similar retrieval-augmented approach for few-shot medical image segmentation has been explored in [34], demonstrating its effectiveness in similar scenarios. In the case of X-ray images, we leverage RAD-DINO [20], a domain-specific version of DINOv2 fine-tuned for radiology, as we expect it to further improve performance.
3 Experiments and results
Datasets. We evaluate In-Context RCA across diverse datasets listed in Figure 2 including SCD [21], HC18 [11], PSFHS [2], PH2 [15], ISIC 2018 [5], NuCLS [1], JSRT [26], 3D-IRCAdB [28], WBC-CV and WBC-JTSC [35]. Unlike previous studies that explored RCA focusing on cardiac MRI [24, 23] or multi-organ segmentation on whole-body MRI [31] only, we broaden our analysis to a wider range of medical imaging challenges. To maintain consistency, we convert all images to grayscale and resize them to 256x256 pixels. For 3D images, we extract slices, keeping those where the region of interest (ROI) is present.
Figure 2: Datasets overview including modality, number of images and structure.
Simulating masks of diverse quality. To evaluate In-Context RCA, we generate segmentation masks of varying quality by training a small U-Net [25] for 10 epochs and saving the lower and higher-quality segmentations. Our goal is to achieve a uniform distribution of mask qualities with varying DSC for a comprehensive evaluation.
Impact of dataset size and RA on performance. We evaluate the impact of the reference dataset size and the use of RA by applying In-Context RCA with UniverSeg as our reverse classifier, considering different support set sizes. We aim to determine if RA maintains or even improves prediction quality while using a smaller reference dataset, thus reducing inference time in practice. We considered reference dataset sizes of 2, 5, 10 and 25 examples. Beyond these values, inference time per image increases substantially. Since UniverSeg supports only binary segmentation, for multi-class we process them separately (including background), and then select the most likely class per pixel using softmax, as described in [4].
Fig. 3 shows the results on HC18 and JSRT datasets, comparing the performance of In-Context RCA for different reference datasets of size $k$. The figure contrasts two selection methods: random selection (first row) and RA (second row), where the $k$ most similar images are chosen based on cosine similarity in the DINOv2 embedding space. The horizontal axis represents the predicted DSC, while the vertical axis shows the real value. As the number of reference samples increases, both methods show improved performance, yielding higher correlation and lower mean absolute error (MAE). However, RA achieves better results with fewer reference samples, what is particularly beneficial during deployment, as it reduces inference time. Other similarity metrics, such as Euclidean distance and inner-product were also considered, but cosine similarity showed the best results.
Figure 3: Scatter plots comparing predicted vs real DSC on HC18 (Left) and JSRT (Right) for different reference datasets of size $k$. Top: results using randomly selected reference samples. Bottom: results following retrieval-augmentation, where the $k$ most similar images are selected based on cosine similarity in the DINOv2 embedding space. In both cases retrieval-augmentation achieves better performance with fewer samples, as shown by higher correlation and lower MAE (bold indicates the best value).
3.1 Comprehensive evaluation across datasets
In order to assess the effectiveness and limitations of In-Context RCA, we conducted an extensive evaluation using UniverSeg and SAM 2 as reverse classifiers, considering DSC as evaluation metric. We compared the performance of In-Context RCA with a single-atlas approach, known to offer the strongest results in the classical RCA framework [31, 24]. For the atlas, we included both traditional and RA versions, whereas for UniverSeg and SAM 2, we focused exclusively on the RA version as it showed better pefromance in our previous experiment (see Fig. 3). We used the segmentation masks generated by a UNet trained for different epochs as previously discussed, to ensure evaluation across varying segmentation qualities. We observe that following RA reference sample selection yields good results without requiring a large reference database.All experiments were conducted on an AMD Ryzen 7 4800H CPU (8 cores) and an NVIDIA GTX 1660 Ti GPU (6GB of VRAM). In all cases, we fixed the size of the reference dataset to 8, based on previous observations.
Fig. 4 provides a comprehensive overview of the results when predicting the DSC across all datasets for all four variants: the traditional RCA framework using single-atlas, its RA version, and the In-Context RCA methods using UniverSeg and SAM 2 as reverse classifiers. We observe that In-Context RCA demonstrates similar and, in some cases, better results compared to the Atlas method, while being significantly more computationally efficient. Particularly in highly irregular segmentation problems like the NuCLS histopathology dataset, our SAM2 variant achieves good correlation of about 0.74 while Atlas methods completely fail. Importantly, while the classic atlas method takes on average one minute to process each image, SAM 2 and UniverSeg achieve the same task in just 0.7 seconds and 0.37 seconds respectively, approximately 85 and 162 times faster. This efficiency, combined with its strong performance, makes In-Context RCA an ideal choice for integration into clinical pipelines.
Figure 4: Scatter plots comparing the predicted DSC across all datasets for In-Context RCA using UniverSeg and SAM 2 as reverse classifiers, followed by the retrieval-augmented and traditional Atlas RCA versions. All evaluations were conducted with a reference dataset of size 8. Bold indicates the winning method per dataset.
4 Conclusion
In this work, we addressed the critical challenge of evaluating medical image segmentation quality without available GT. To do so we introduced In-Context RCA, a novel framework that builds on top of RCA, leveraging the power of in-context segmentation models as reverse classifiers. By also incorporating retrieval augmentation, our method enables efficient quality estimation using smaller reference databases while either maintaining or improving prediction accuracy.
Experimental results conducted across diverse medical imaging modalities, show that In-Context RCA achieves comparable or superior performance to traditional atlas-based approaches while being significantly more computationally efficient. It also overcomes the limitations imposed by atlas-based methods allowing to tackle scenarios beyond anatomical segmentation, such as cell images. Overall, our results position In-Context RCA as a promising solution for automated quality control in clinical workflows, where rapid assessment is essential.